Let A and B be events such that P(A) = 0.6, P(A È B) = 0.8 and P(A | B) = 0.6.

Find P(B).



1
Expert's answer
2021-10-06T12:14:03-0400

"P(A\ and\ B)=P(A)P(B|A)=0.6\cdot 0.8=0.48"


"P(B)=\frac{P(A\ and\ B)}{P(A|B)}=\frac{0.48}{0.6}=0.8"